OpenStudy (anonymous):

what is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form.

OpenStudy (ranga):

Given two points do you know how to find the slope?

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1) now apply (y-y1)=slope(x-x1) slope comes out to be 4 apply (y-1)=4(x-4) y-1=4x-16 y+15=4x is the eqn of the reqd line